# Cascade Speculative Drafting for Even Faster LLM Inference

### 21 Dec 2023 v2

@@ Abstract
Speculative decoding reduces the latency of large language model (LLM) inference by leveraging a draft model to draft for a larger target model to review. However, drafting in speculative decoding involves slow autoregressive generation and allocates the same amount of time to generating tokens of varying importance. These two inefficiencies lead to its suboptimal performance. To address this issue, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade constitutes efficient time allocation in drafting. Combining both cascades, our CS Drafting algorithm has achieved up to a 72 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
@@ 1 Introduction
The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but have also found widespread application in user-interactive settings such as chatbots and virtual assistants. This application involves an extremely high number of users, up to hundreds of millions daily. To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the provider and the community. Consequently, optimizing LLM latency has become a crucial and impactful research focus.
@@
Unfortunately, LLMs currently suffer from high latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel. The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself. As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two main factors: the expected number of tokens accepted by the target model per run, and the latency of drafting.
@@
Figure 1: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models. The draft model generates 30 tokens at each step.
Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens. The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens proposed to and verified by LLMs at the next outer layer, $\mathcal{M}_{d(i+1)}$. This recursive process forms the vertical cascade. After a vertical cascade, a smaller $\mathcal{M}_{d(i)}$ will serve as the target model and perform a vertical cascade. The process repeats until the smaller model becomes a statistical language model that will generate autoregressively forming the horizontal cascade.
[…]
In light of this, there exist a type of draft models, including statistical language models like bigram language models, that are immune to drafting efficiency issues. Owing to their simple structure, these models incur negligible latency and computational resource costs compared to neural language models, thereby avoiding the aforementioned trade-off. Therefore, assuming enough parallel computing power to perform computation on additional $k$ tokens with similar time cost (Leviathan et al., 2023), utilizing speculative decoding with a statistical language model as the draft model can generally and intuitively enhance the efficiency of a neural language model. However, due to their limited capabilities, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the target model when the neural language model is available.
Based on the above analysis, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model. Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model. We refer to this tiered speculative decoding approach as the Vertical Cascade.
@@
Furthermore, our analysis in Figure 1 reveals an interesting pattern in the expected acceptance rate during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model [^1] A theoretical analysis is provided in Section 4. Note that this probability differs from the acceptance rate, which is the probability of a token being accepted conditioned on the previous tokens’ acceptance.. This indicates that later tokens from draft models, more prone to rejection, contribute less to the expected number of accepted tokens, yet incur the same latency. Based on this observation and our practice, we suggest the use of smaller, faster draft models for generating these high-rejection tokens, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 2.
@@
Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 72% over speculative decoding. These findings highlight the practical advantages and efficiency enhancements offered by both vertical and horizontal cascades.
Our main contributions are summarized as follows:
@@ 2 Preliminary
2.1 Speculative Decoding Algorithm
In real-world applications of large language models (LLMs), the sequential nature of autoregressive decoding emerges as a primary latency bottleneck in terms of response time. Addressing this, speculative decoding (SD) (Leviathan et al., 2023) presents an innovative solution that utilizes a compact draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$ In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$. At the end, the target model will generate one additional token following the accepted tokens. Such a design guarantees the output is the same as sampling autoregressively using the target model alone (Leviathan et al., 2023).
This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of LLMs.
3 Cascaded Speculative Drafting
Algorithm 1 CascadeSpeculativeDrafting
  Require:draft model $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target model $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
@@ 3 Cascaded Speculative Drafting
   $\triangleright$ Initialize curGen and curProb
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones of length prefix
   $\triangleright$ Unpack the a list of $k$ for current function call
@@
   $\triangleright$ Base case of the recursive call when the MaG generate
  if draftList only contains one element then
@@
      $\triangleright$ MaG generates greedily, so the logits will be 1 at the position of the generation token and 0 otherwise
@@
   $\triangleright$ Begin the cascade
@@
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,j)$ extending to the bottom-right corner
     while length of curGen - prefix.length is less than $k_{i}$ do
        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, False, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{r}]$
         $\triangleright$ Do not overwrite the previous probabilities
[…]
        Add elements of $[p_{s},p_{2},...,p_{r}]$ to curProbs
@@
   $\triangleright$ No leniency when the target model reviews
@@
   $\triangleright$ Review process by $\mathcal{M}_{t}$ to ensure the output distribution is the same as autoregressive generation
  $[p_{1}^{\prime},p_{2}^{\prime},...,p_{m}^{\prime}]\leftarrow$ $\mathcal{M}_{t}(curGen)$
  $index\leftarrow m$
  for $i\leftarrow 1+prefix.length$ to $m-1$ do
     Sample $r$ from a uniform distribution
     if $r\leq min(1,\frac{lp_{i}^{\prime}}{p_{i}})$ then
        $index\leftarrow i$ and exit the loop
     end if
  end for
  Sample $x_{last}$ from $p_{index}^{\prime}-p_{index}$
  return $[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
@@ 3.1 Vertical Cascade
The speculative decoding algorithm relies on the autoregressive generation of a smaller draft model. However, despite its smaller size, the draft model must run $k$ times for each target model run, and the cost can still be significant. To further reduce latency in the drafting process, we propose using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model. This process can be done recursively until it reaches the smallest draft model. The smallest draft model can be a statistical language model, such as a bigram language model, where the cost is negligible compared to neural language models. We refer to this recursive process as Vertical Cascade.
Furthermore, to further improve speed, we can introduce leniency, a hyperparameter in speculative decoding that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023). With leniency $l\in[1,\infty)$, the acceptance condition for token $x_{i}$ in speculative decoding becomes $\mathcal{M}{d}(x)[i]<l\times\mathcal{M}{t}(x)[i]$. Otherwise, with a probability of $1-\frac{l\times\mathcal{M}{t}(x)}{\mathcal{M}{d}(x)}$, we reject $x_{i}$ and any following tokens. While the reduced quality introduced by leniency in speculative decoding is generally undesirable in practice, it fortunately does not affect the final output in our vertical cascade approach—we can only apply leniency in the inner speculative decoding. As long as leniency is not applied to the target model, the final quality is always guaranteed by the target model.
@@ 3.2 Horizontal Cascade
As illustrated in Figure 1, another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal. The first draft token is more likely to be accepted as it only depends on itself; the last token is rarely accepted, as it has a chance of being reviewed only if all preceding tokens are accepted. From another perspective, one could also say that the first draft token bears the responsibility for all succeeding tokens, while the last token bears none. In Section 4, we provide a theoretical analysis of this phenomenon.
Based on this observation, we propose better allocation of the time cost of drafting by using larger models to generate the preceding draft tokens and smaller models for the succeeding tokens. This design allows more time to be allocated to more effective tokens, leading to a reduction in overall latency. We refer to this design as the Horizontal Cascade.

### 16 Feb 2024 v3

@@ Abstract
Introduced to enhance the efficiency of large language model (LLM) inference, speculative decoding operates by having a smaller model generate a draft. A larger target model then reviews this draft to align with its output, and any acceptance by the target model results in a reduction of the number of the target model runs, ultimately improving efficiency. However, the drafting process in speculative decoding includes slow autoregressive generation and allocates equal time to generating tokens, irrespective of their importance. These inefficiencies collectively contribute to the suboptimal performance of speculative decoding. To further improve LLM inference, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade optimizes time allocation in drafting for improved efficiency. Combining both cascades, CS Drafting achieves up to an 81 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model. Our code is publicly available at https://github.com/lfsszd/CS-Drafting.
@@ 1 Introduction
The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023), has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but have also found widespread applications in user-interactive settings, such as chatbots and virtual assistants. However, these applications involve an extremely high number of users, up to hundreds of millions daily. To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the service provider and the community. Consequently, optimizing the latency of LLMs has become a critical area of research.
@@
Unfortunately, the ever-growing size of LLMs significantly increases the latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter) autoregressively, and these tokens are then reviewed by a target model (which is larger and slower) in parallel. In one single run, the target model will accept any tokens aligned with its output and further generate one token. The drafting process in speculative decoding enables the target model to generate multiple tokens in a single run while maintaining its output distribution unchanged. With a properly sized draft model, speculative decoding achieves a speedup of 2 to 3 times, making it a potential method for solving high latency issues.
Figure 1: The figure shows an example of Cascade Speculative Drafting with target model $\mathcal{M}_{t}$ and draft models $\mathcal{M}_{d_{1}}$, $\mathcal{M}_{d_{2}}$, and $\mathcal{M}_{d_{3}}$. The horizontal cascade involves using larger draft models to generate the earlier tokens and smaller models to generate the later tokens. The vertical cascade requires each model to review drafts from smaller models with the exception of the smallest model which is a statistical language model. As the horizontal cascade and vertical cascade are orthogonal to each other, CS Drafting combines both approaches for optimal efficiency.
However, since draft models are typically required to generate multiple tokens in multiple steps, where each generation still involves inefficient autoregressive decoding, the performance of speculative decoding could be limited by the drafting latency. This inefficiency is also indicated by Leviathan et al. (2023)’s experiments, where it was observed that very small models (e.g., around two orders of magnitude smaller than the target model) are usually the best choice for drafting because their inference cost is lower compared to that of a larger draft model, despite the fact that larger draft models usually have higher-quality generation. This underscores that improving drafting efficiency is crucial for further enhancing the performance of speculative decoding.
In light of this, one key strategy to address this bottleneck is to avoid the inefficient autoregressive generation of neural draft models. Based on this consideration, it is noted that statistical language models, such as bigram language models, incur negligible latency and computational resource costs compared to neural language models, owing to their simple structure. However, because the tokens generated by statistical language models usually do not have a high probability of being accepted by the target model, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the draft model. Nonetheless, we notice that it is not necessary to use only one draft model in speculative decoding—statistical language models can serve as the “draft” model for the neural draft model, thereby eliminating autoregressive generation from the neural draft model.
@@
Figure 2: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on GSM8K. The draft model generates 30 tokens at each step. The results on MMLU are shown in Figure 3 of the Appendix.
@@
Furthermore, our analysis in Figure 2 reveals a pattern during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model. This is because the probability of a token being accepted is conditioned on the acceptance of the previous tokens. It indicates that later tokens from draft models are more prone to rejection, contributing less to the expected number of accepted tokens per draft step, yet incurring the same latency.
Inspired by the above observations, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model. Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model. In this design, the drafting of each neural model will be accelerated by drafting from a smaller model, avoiding the inefficiency of autoregressive generation from neural models. We refer to this tiered speculative decoding approach as the Vertical Cascade. In addition, we suggest the use of smaller, faster draft models for generating high-rejection tokens that are trailing in drafting generation, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 1.
@@
Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 81% over speculative decoding. These findings highlight the practical advantages and efficiency enhancements offered by both vertical and horizontal cascades.
The main contributions are summarized as follows:
@@ 2 Preliminary
The core concept of speculative decoding (Leviathan et al., 2023) involves the utilization of a small draft model for sequential token generation with validation by a larger target model resulting in reduced latency. This design accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, $k$ be the number of draft tokens generated per step, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$. In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]\leq\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $norm(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$. At the end, the target model will generate one additional token following the accepted tokens. Such a design guarantees the output is the same as sampling autoregressively using the target model alone, as proven in Leviathan et al. (2023).
Speculative decoding was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for reducing the latency of LLM inference.
3 Cascade Speculative Drafting
In this section, we introduce our proposed method, Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades: vertical cascade and horizontal cascade.
Algorithm 1 CascadeSpeculativeDraftingStep
  Require: draft models $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target mode $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
@@ 3 Cascade Speculative Drafting
   $\triangleright$ Initialize curGen and curProb.
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones with the same length as prefix
   $\triangleright$ Unpack the a list of $k$ for the current function call.
@@
   $\triangleright$ Generate using MaG for the Base case of the recursive call.
  if draftList is empty then
@@
   $\triangleright$ Perform the horizontal cascade with the for loop.
@@
      $\triangleright$ Prepare the arguments for the next recursive call.
@@
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,i+1)$ extending to the bottom-right corner curPrefix $\leftarrow$ curGen
     while curGen.length - curPrefix.length is less than $k_{i}$ do
        curPrefix $\leftarrow$ curGen
         $\triangleright$ Perform the vertical cascade with the recursive call.
        $[x_{1},..,x_{u}],[p_{1},p_{2},...,p_{v}]\leftarrow$ CascadeSpeculativeDraftingStep(curDraftList, curTarget, curPrefix, False, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{u}]$
        s $\leftarrow$ curProbs.length + 1
        Add elements of $[p_{1},p_{2},...,p_{v}]$ to curProbs
@@
   $\triangleright$ Set lenience to $1$ when the original target model reviews.
@@
   $\triangleright$ Use $\mathcal{M}_{t}$ to review the draft generation.
  $[x_{1},...,x_{out}],[p_{1}^{\prime},p_{2}^{\prime},...,p_{out}^{\prime}]$ = review($\mathcal{M}_{t}$, curGen, curProbs, l)
  return $[x_{1},...,x_{out}],[p_{1}^{\prime},p_{2}^{\prime},...,p_{out}^{\prime}]$
@@ 3.1 Vertical Cascade
A notable inefficiency of the speculative decoding algorithm is the reliance on the autoregressive generation of a smaller draft model. Since the draft model must run $k$ times for each target model run, the cost can still be significant despite its smaller size. In light of this, we reduce the drafting inefficiency by using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model. In addition, since this process can be performed again on the draft model that drafts for the original model, we recursively perform this process until it reaches a statistical draft model that involves negligent cost, such as a bigram language model. In this approach, we expect each recursion step will reduce the drafting latency without altering the output distribution. We refer to this recursive speculative approach as Vertical Cascade.
Additionally, we incorporate lenience, a hyperparameter that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023). Lenience can be adopted during sampling or greedy decoding with speculative decoding. Let lenience $l\in[1,\infty)$. When sampling, the acceptance condition for token $x_{i}$ is transformed to $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$. If the acceptance condition is not satisfied, with a probability of $1-\frac{l\times\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$, we reject $x_{i}$ and any following tokens. [^1] For simplicity, we assume the probability outputs are not standardized. We refer the readers to Leviathan et al. (2023) for the discussion on standardized sampling. When performing greedy decoding, the acceptance condition becomes deterministic and is simply either $argmax\mathcal{M}_{d}(x)[i]=argmax\mathcal{M}_{t}(x)[i]$ or $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$.
For the speculative decoding algorithm, the reduced quality introduced by lenience is generally undesirable. However, for the vertical cascade approach, lenience affects the final output only if it is applied when the target model reviews. Therefore, we can limit the application of lenience in the vertical cascade only when draft models review and do not apply lenience when the target model reviews. This can ensure the final output is not altered while further reducing latency.
@@ 3.2 Horizontal Cascade
Another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal, as illustrated in Figure 2. The first draft token is more likely to be accepted as it only depends on itself; the last token is rarely accepted, as it has a chance of being reviewed only if all preceding tokens are accepted. From a theoretical perspective, assume the event of acceptance of each token being a Bernoulli distribution with probably $p$, the probability of $n$-th token being accepted is $p^{n}$, implying an exponential decrease of value for tokens generated later in the sequence.
Inspired by this observation, we designed Horizontal Cascade, an approach that improves time allocation by draft token allocation. Horizontal Cascade assigns the largest draft model to perform the generation of the first draft token due to its highest output alignment with the target model, and it progressively uses a smaller as the new draft token to be generated is less likely to be accepted. This process stops after the smallest model, i.e., a statistical language model finishes. This design reduces the time cost of generating unimportant draft tokens with a costly draft model, leading to a reduction in overall latency.